# Exploring the interaction between the MW and LMC with a large sample of blue horizontal branch stars from the DESI survey

### v1

@@ Abstract
pubyear: 2024pagerange: Exploring the interaction between the MW and LMC with a large sample of blue horizontal branch stars from the DESI surveyD
@@ 2.1 The DESI spectroscopic survey
DESI began its 5 month long survey validation campaign in December 2020 (DESI Collaboration et al., 2024d), and the scientific part of the survey began in May 2021. In June 2023, DESI published its early data release (EDR) [^2] https://data.desi.lbl.gov/doc/releases/edr/, which consists of the survey validation data (DESI Collaboration et al., 2024e). The MWS published a stellar value-added catalogue based on this EDR that contains radial velocity and stellar parameter measurements for about 400,000 unique stars (Koposov et al., 2024). The upcoming data release, DR1, will consist of data observed until June 2022, and is expected by spring of 2025. Since then, the survey has progressed and has as of writing completed 70  per cent of observations in dark conditions and 75  per cent of observations in bright conditions.
The work in this paper is based on the upcoming DR1, as well as spectra observed in the second year of DESI observations between June 2022 and June 2023 that are not yet prepared for public release. This latter data will be referred to as year 2 data, or Y2. The DR1 dataset consists of coadditions (or coadds), i.e. averages, of spectra processed similarly to the EDR (DESI Collaboration et al., 2024e). In the EDR, spectra of objects in a given survey and program are coadded within HEALPixels on the sky using an $N_{\text{side}}=64$ nested HEALPix tesselation. One HEALPixel can cover several tiles so that spectra of objects that were observed in different tiles are still coadded (Guy et al., 2023). The Y2 dataset is different, because it has been based on daily processing of DESI data and therefore consists of stacked spectra within individual DESI tiles (Schlafly et al., 2023). This means that stars that were observed on multiple tiles will not have their spectra coadded. In short, DR1 data is coadded across HEALPixels and Y2 across tiles.
Using DR1 (DESI Collaboration et al 2024 I, in prep.), the DESI collaboration has released so-called key papers presenting the two point clustering measurements and validation (DESI Collaboration et al 2024 II, in prep.), baryonic acoustic oscillations measurements from galaxies and quasars (DESI Collaboration et al., 2024a) and from the Lyman $\alpha$ forest (DESI Collaboration et al., 2024b), as well as a full-shape analysis of galaxies and quasars (DESI Collaboration et al 2024 III, in prep.). There are cosmological results from the baryonic acoustic oscillations measurements (DESI Collaboration et al., 2024c) and the full-shape analysis (DESI Collaboration et al 2024 IV, in prep.), as well as constraints on primordial non-gaussianities (DESI Collaboration et al 2024 V, in prep.)
@@ 2.3 BHB stars in the MWS
The MWS contains a high-priority sample. It has higher priority than the main and faint samples, which in order of highest priority to lowest is a sample of white dwarfs, RRLe, nearby stars, and BHB stars. These targets are prioritised when DESI assigns fibers for observations, to ensure high completeness. There are in total 10,695 BHB targets to be included in DR1 and that thus are coadded by HEALPixel, and 10,498 in Y2, coadded by tiles. Both within DR1 and within Y2, and in between DR1 and Y2, there are repeat observations of the same star, so these duplicates will be removed. DR1 and Y2 BHBs also includes both main survey and secondary targets, and we will introduce them separately below.
@@
where $G$ is the Gaia parallax, $\varpi$ is the Gaia parallax and $\sigma_{\varpi}$ the associated error, $g$ and $r$ are DESI Legacy Imaging Survey optical photometry, and
$X_{\text{BHB}}=$ $(g-z)-[1.0716(g-r)^{5}-1.42272(g-r)^{4}$ (1)
@@
which is a fit to the BHB sequence on the CMD, and
@@ 2.4 Cleaning the BHB sample
BSs are stars that appear more massive and younger than stars of the same populations and thus appear bluer than stars of the same luminosity, because they formed through mergers or mass-transfer between two or three stars (Perets & Fabrycky, 2009). BS have similar colours to BHBs, and so they contaminate simple BHB colour-based selections, as e.g. the target selection for BHBs. Eq. (1) for the main survey BHB targets and the requirements on both $(g-r)$, $(r-z)$ for the secondary BHB targets are designed to remove most of this BS contamination. However to not compromise BHB completeness, these colour cuts are not very aggressive, so we wish to remove BS stars without removing more BHBs. These remaining BS can be removed by utilising spectroscopic information derived from the RVS pipeline, see e.g. Clewley et al. (2002); Xue et al. (2008).
@@ 2.5 BHB distances
The DESI BHB sample contains a large number of Sgr stream stars, which has been mapped with RRLe by Hernitschek et al. (2017). They provide a spline for heliocentric distances and Sgr stream longitude $\tilde{\Lambda}$ for the leading arm, that we interpolate to create a function of heliocentric distances as a function of $\tilde{\Lambda}$ (depicted in orange in Fig. 3). In Sec. 2.6 we introduce the selection of Sgr stream members using this interpolated spline, and we apply this selection here as well to identify Sgr leading arm stars. Then for all stars that we flag as belonging to the Sgr leading arm in the range $50^{\circ}<$ $\tilde{\Lambda}$ $<100^{\circ}$, we compute the distances derived using the relationship in BK15 $d_{\mathrm{BK15}}$, and divide them by the RRLe-inferred Sgr distances $d_{\mathrm{Sgr}}$. The fraction $d_{\mathrm{BK15}}/d_{\mathrm{Sgr}}$ is 1.080. This means that the BHB distances derived by using the relationship in BK15 are overestimated by around 8  per cent, where we assume that the RRLe-derived distances are correct. A part of this uncertainty might also come from the RRLe distances, but nevertheless, the colour-magnitude relation must be corrected. The 8  per cent overestimation implies a distance modulus shift of $-5\log_{10}(1.080)=-0.167$; to undo this shift, this value is subtracted from the BK15 formula and its linear extrapolation to compute $M_{g,\mathrm{Sgr-corr}}$. The final BHB absolute magnitude-colour relation we get is:
@@
We validate these distances against GCs next. We do this by taking stars from Vasiliev & Baumgardt (2021) with more than 90  per cent probability of being GC members, which were selected based on Gaia proper motions and parallaxes, and crossmatch with DECaLS DR9 to get their dereddened photometry. BHB stars are photometrically selected as GC members with $-0.3<(g-r)<0$ and $-0.15<(r-z)-(g-r)<0.02$, which yields 803 BHBs in 23 GCs. Using the GC literature distances from Vasiliev & Baumgardt (2021) for each star we compute its absolute magnitude $M_{g,\mathrm{GC}}$, and computing the absolute magnitude $M_{g,\mathrm{thiswork}}$ using the DECaLS magnitudes in Eq. (6), we can compute the distance modulus offset $\Delta M_{g}=M_{g,\mathrm{thiswork}}-M_{g,\mathrm{GC}}$. This distribution shows that the photometric selection contains outliers that likely are BSs, so we make an additional cut in $|\Delta M_{g}|<1$, which yields 722 BHBs in 22 GCs. We find that the median of $\Delta M_{g}$ is $0.093$ mag, which implies a systematic distance underestimation of about 4.2  per cent, and the standard deviation is 0.133 mag. This systematic offset is likely caused by a metallicity dependence in absolute magnitudes of BHBs (Fermani & Schönrich, 2013), which will be explored further in future work.
@@ 2.6 Removing halo substructure
The BHB sample before and after the Sgr removal can be seen to the left and right in Fig. 3 respectively, with the right-hand side of the figure clearly showing the Sgr overdensities disappearing after the selection has been applied. The top row shows the stars’ heliocentric distances as a function of $\tilde{\Lambda}$; it is in this space, combined with the third axis of $\tilde{B}$, that the Sgr selection is made. The bottom row shows the radial velocity in Galactic standard of rest (GSR), $v_{\text{GSR}}$ [^6] https://docs.astropy.org/en/stable/generated/examples/coordinates/rv-to-gsr.html was used for computing $v_{\text{GSR}}$., as a function of $\tilde{\Lambda}$, which is an independent check that the removal of Sgr stars worked: Sgr in our sample clearly aligns with the Sgr velocity model from Vasiliev et al. (2020), and even though we make no explicit selections in velocity space, the stars on Sgr-like velocities are also removed from the sample. In the upper left panel at $\tilde{\Lambda}$ $\sim 60-100$ degrees, we see that the Sgr leading arm appears vertically extended, which can be explained by the bifurcation of the stream leading arm (Belokurov et al., 2006). The top right of the same figure shows the Sgr leading arm apocentre (($\tilde{\Lambda}$, distance) $\sim(70^{\circ}$, 60 kpc)). This apocentre has been observed in RRLe (Sesar et al., 2017b; Hernitschek et al., 2017) and simulations (Gibbons et al., 2014) to contain a spur-like structure that is also present in our data and removed with our Sgr removal defined above.
@@ 3.1 Maximum likelihood estimation
To characterise the velocity distribution of the halo, we need to analyse the average motions of the BHBs selected by their spatial positions. We will model the stars by assuming that their velocities can be described by a Gaussian distribution with a mean $\mu$ and a standard deviation $\sigma$. When evaluating the likelihood for each star, we also take individual uncertainties on the radial velocity measurements $\epsilon_{i}$, for each observed velocity $v_{r,i}$, into account. We need to find the underlying velocity distribution that maximizes the likelihood of observing the measured velocities. Thus, characterising the halo’s velocity field is a maximum likelihood estimation (MLE) problem. The log-likelihood that we want to maximize is:
@@ 3.3 Parametrizing the halo velocity field
We can find $v_{\text{compr}}$, $v_{\text{dipole}}$, $l_{\text{apex}}$ and $b_{\text{apex}}$ by once again using MLE: we want to find the four parameter values (and a dispersion $\sigma$) that maximizes the likelihood of observing the velocities of our outer halo BHB sample. We do this by replacing $\mu$ in Eq. (8) with $v_{\text{field}}$, keeping the observed velocities $v_{r}$ and their corresponding errors $\epsilon_{i}$ the same as in the previous numerical maximization. Instead of using data in Voronoi cells on the sky as we did previously, we now fit the entire data set. Because of the geometry of the sphere, we require that $b_{\text{apex}}$ lies in the range $-90^{\circ}$ to $90^{\circ}$ as allowing the latitude to extend past these values is unphysical, but impose no bounds on the other parameters to remain agnostic. The parameter errors are again computed using Eq. (9).
@@
The same parameters as presented in Table 1 are also derived by PP21 and Y24, and so we can compare their results to ours [^8] Both works use several different tracer populations, but we only compare to the results using their BHB sample.. We can interpret our $v_{\text{compr}}$ and $v_{\text{dipole}}$ as being the same parameter as their mean halo radial velocities and disc barycentric velocities. These, as well as our $b_{\text{apex}}$, are consistent between all papers using BHB stars as tracers. However, our $l_{\text{apex}}$ does not agree with either work; this means that our velocity field has a different direction than their studies. The resulting the angular offsets between the apex directions are $76^{\circ}$ and $57^{\circ}$, respectively. This substantial difference means that the velocity field models predict different velocities across the sky. Our apex direction ($l_{\text{apex}},b_{\text{apex}}$) is seen on-sky in Fig. 9 in blue together with the BHB sample in light grey, and the apex directions found by Y24, PP21 and Chandra et al. (2024) in pink, magenta and purple, respectively, with the position of the LMC as an orange star with its past orbit also in orange.
@@
Figure 9: The apex direction longitude and latitude ($l_{\text{apex}},b_{\text{apex}}$) from Table 1 shown on-sky in blue, and the apex directions from Chandra et al. (2024), PP21 and Y24 as a purple square, magenta upwards triangle and a violet downwards triangle respectively. For PP21 we use their BHB-derived value, and for Y24, we use their BHB-derived apex direction for the distance bin with Galactocentric distances larger than 50 kpc. The distribution of the BHB sample is shown in light gray. The position of the LMC is shown as an orange star with its past trajectory in orange (Vasiliev, 2023).

### v3

@@ Abstract
pubyear: 2025pagerange: Exploring the interaction between the MW and LMC with a large sample of blue horizontal branch stars from the DESI surveyE
@@ 2.1 The DESI spectroscopic survey
DESI began its 5 month long survey validation campaign in December 2020 (DESI Collaboration et al., 2024b), and the scientific part of the survey began in May 2021. In June 2023, DESI published its early data release (EDR) [^2] https://data.desi.lbl.gov/doc/releases/edr/, which consists of the survey validation data (DESI Collaboration et al., 2024c). The MWS published a stellar value-added catalogue based on this EDR that contains radial velocity and stellar parameter measurements for about 400,000 unique stars (Koposov et al., 2024). The first data release (DR1) [^3] https://data.desi.lbl.gov/doc/releases/dr1/ was published in March 2025 and contains observations taken between May 2021 through June 2022 (DESI Collaboration et al., 2025a), with a value-added catalogue published by the MWS containing over 6 million sources with radial velocities and stellar parameters (Koposov et al., 2025). The BHB sample produced from DR1 data in this work was also made public as a value-added catalogue [^4] https://data.desi.lbl.gov/doc/releases/dr1/vac/mws-bhb/. Since then, the survey has progressed and has as of writing completed 93  per cent of observations in dark conditions and 96  per cent of observations in bright conditions.
The work in this paper is based on DR1 (DESI Collaboration et al., 2025a), as well as spectra observed in the second year of DESI observations between June 2022 and June 2023 that are not yet prepared for public release. This latter data will be referred to as year 2 data, or Y2. The DR1 dataset consists of coadditions (or coadds), i.e. averages, of spectra processed similarly to the EDR (DESI Collaboration et al., 2024c). In the EDR, spectra of objects in a given survey and program are coadded within HEALPixels on the sky using an $N_{\text{side}}=64$ nested HEALPix tesselation. One HEALPixel can cover several tiles so that spectra of objects that were observed in different tiles are still coadded (Guy et al., 2023). The Y2 dataset is different, because it has been based on daily processing of DESI data and therefore consists of stacked spectra within individual DESI tiles (Schlafly et al., 2023). This means that stars that were observed on multiple tiles will not have their spectra coadded. In short, DR1 data is coadded across HEALPixels and Y2 across tiles.
Using DR1 (DESI Collaboration et al., 2025a), the DESI collaboration has released so-called key papers presenting the two point clustering measurements and validation (DESI Collaboration et al., 2025e), baryonic acoustic oscillations measurements from galaxies and quasars (DESI Collaboration et al., 2025d) and from the Lyman $\alpha$ forest (DESI Collaboration et al., 2025b), as well as a full-shape analysis of galaxies and quasars (DESI Collaboration et al., 2024a). There are cosmological results from the baryonic acoustic oscillations measurements (DESI Collaboration et al., 2025c) and the full-shape analysis (DESI Collaboration et al., 2025f).
@@ 2.3 BHB stars in the MWS
The MWS contains a high-priority sample. It has higher priority than the main and faint samples, which in order of highest priority to lowest is a sample of white dwarfs, RRLe, nearby stars, and BHB stars. These targets are prioritised when DESI assigns fibers for observations, to ensure high completeness. There are in total 10,695 BHB targets included in DR1 and that thus are coadded by HEALPixel, and 10,498 in Y2, coadded by tiles. Both within DR1 and within Y2, and in between DR1 and Y2, there are repeat observations of the same star, so these duplicates will be removed. DR1 and Y2 BHBs also include both main survey and secondary targets, and we will introduce them separately below.
@@
where $G$ is the Gaia magnitude, $\varpi$ is the Gaia parallax and $\sigma_{\varpi}$ the associated error, $g$ and $r$ are DESI Legacy Imaging Survey optical photometry, and
$X_{\text{BHB}}$ $=(r-z)-[1.07163(g-r)^{5}-1.42272(g-r)^{4}$ (1)
@@
which is a boundary of BHBs and BSs in colour-colour space that is used to remove BS contamination from the BHB sample [^7] Note that there is a typo in Eq. (1) in Cooper et al. (2023) where the first term of the equation is $(g-z)$, which is supposed to be $(r-z)$ as in this work. (this relation is originally from the $S^{5}$ survey described in Li et al. (2019) and adopted for the MWS, see their Eq. (6) and the corresponding section for more details), and
@@ 2.4 Cleaning the BHB sample
BSs are stars that appear more massive and younger than stars of the same populations and thus appear bluer than stars of the same luminosity, because they formed through mergers or mass-transfer between two or three stars (Perets & Fabrycky, 2009). BS have similar colours to BHBs, and so they contaminate simple BHB colour-based selections, as e.g. the MWS target selection for BHBs. Eq. (1) for the main survey BHB targets and the requirements on both $(g-r)$, $(r-z)$ for the secondary BHB targets are designed to remove most of this BS contamination. However to not compromise BHB completeness, these colour cuts are not very aggressive, so we wish to remove BS stars without removing more BHBs. These remaining BS can be removed by utilising spectroscopic information derived from the RVS pipeline, see e.g. Clewley et al. (2002); Xue et al. (2008).
@@ 2.5 BHB distances
The DESI BHB sample contains a large number of Sgr stream stars, which has been mapped with RRLe by Hernitschek et al. (2017). They provide a spline for heliocentric distances and Sgr stream longitude $\tilde{\Lambda}$ for the leading arm, that we interpolate to create a function of heliocentric distances as a function of $\tilde{\Lambda}$ (depicted in orange in Fig. 3). In Sec. 2.6 we introduce the selection of Sgr stream members using this interpolated spline, and we apply this selection here as well to identify Sgr leading arm stars. Then for all stars that we flag as belonging to the Sgr leading arm in the range $50^{\circ}<$ $\tilde{\Lambda}$ $<100^{\circ}$, we compute the distances derived using the relationship in BK15 $d_{\mathrm{BK15}}$, and divide them by the RRLe-inferred Sgr distances $d_{\mathrm{Sgr}}$. The fraction $d_{\mathrm{BK15}}/d_{\mathrm{Sgr}}$ is 1.080. This means that the BHB distances derived by using the relationship in BK15 are overestimated by around 8  per cent, where we assume that the RRLe-derived distances are correct. A part of this uncertainty might also come from the RRLe distances, but nevertheless, the colour-magnitude relation must be corrected. The 8  per cent overestimation implies a distance modulus shift of $-5\log_{10}(1.080)=-0.167$; to undo this shift, this value is subtracted from the BK15 formula and its linear extrapolation to compute $M_{g}$. The final BHB absolute magnitude-colour relation we get is,
@@
We validate these distances against GCs next. We do this by taking stars from Vasiliev & Baumgardt (2021) with more than 90  per cent probability of being GC members, which were selected based on Gaia proper motions and parallaxes, and crossmatch with DECaLS DR9 to get their dereddened photometry. BHB stars are photometrically selected as GC members with $-0.3<(g-r)<0$ and $-0.15<(r-z)-(g-r)<0.02$, which yields 803 BHBs in 23 GCs. Using the GC literature distances from Vasiliev & Baumgardt (2021) for each star we compute its absolute magnitude $M_{g,\mathrm{GC}}$, and computing the absolute magnitude $M_{g}$ using the DECaLS magnitudes in Eq. (6), we can compute the distance modulus offset $\Delta M_{g}=M_{g}-M_{g,\mathrm{GC}}$. This distribution shows that the photometric selection contains outliers that likely are BSs, so we make an additional cut in $|\Delta M_{g}|<1$, which yields 722 BHBs in 22 GCs. We find that the median of $\Delta M_{g}$ is $0.093$ mag, which implies a systematic distance underestimation of about 4.2  per cent, and the standard deviation is 0.133 mag. This systematic offset is likely caused by a metallicity dependence in absolute magnitudes of BHBs (Fermani & Schönrich, 2013), which will be explored further in future work.
@@ 2.6 Removing halo substructure
The BHB sample before and after the Sgr removal can be seen to the left and right in Fig. 3 respectively, with the right-hand side of the figure clearly showing the Sgr overdensities disappearing after the selection has been applied. The top row shows the stars’ heliocentric distances as a function of $\tilde{\Lambda}$; it is in this space, combined with the third axis of $\tilde{B}$, that the Sgr selection is made. The bottom row shows the radial velocity in Galactic standard of rest (GSR), $v_{\text{GSR}}$ [^9] https://docs.astropy.org/en/latest/coordinates/example_gallery_rv_to_gsr.html was used for computing $v_{\text{GSR}}$., as a function of $\tilde{\Lambda}$, which is an independent check that the removal of Sgr stars worked: Sgr in our sample clearly aligns with the Sgr velocity model from Vasiliev et al. (2020), and even though we make no explicit selections in velocity space, the stars on Sgr-like velocities are also removed from the sample. In the upper left panel at $\tilde{\Lambda}$ $\sim 60-100$ degrees, we see that the Sgr leading arm appears vertically extended, which can be explained by the bifurcation of the stream leading arm (Belokurov et al., 2006). The top right of the same figure shows the Sgr leading arm apocentre (($\tilde{\Lambda}$, distance) $\sim(70^{\circ}$, 60 kpc)). This apocentre has been observed in RRLe (Sesar et al., 2017b; Hernitschek et al., 2017) and simulations (Gibbons et al., 2014) to contain a spur-like structure that is also present in our data and removed with our Sgr removal defined above.
@@ 3.1 Maximum likelihood estimation
To characterise the velocity distribution of the halo, we need to analyse the average motions of the BHBs selected by their spatial positions. We will model the stars by assuming that their velocities can be described by a Gaussian distribution with a mean $\mu$ and a standard deviation $\sigma$. When evaluating the likelihood for each star, we also take individual uncertainties on the radial velocity measurements $\epsilon_{i}$, for each observed velocity $v_{r,i}$, into account. We need to find the underlying velocity distribution that maximizes the likelihood of observing the measured velocities. Thus, we characterise the halo’s velocity field using maximum likelihood estimation (MLE). The log-likelihood that we want to maximize is:
@@ 3.3 Parametrizing the halo velocity field
We can find $v_{\text{compr}}$, $v_{\text{dipole}}$, $l_{\text{apex}}$ and $b_{\text{apex}}$ by once again using MLE: we want to find the four parameter values (and a dispersion $\sigma$) that maximizes the likelihood of observing the velocities of our outer halo BHB sample. We do this by replacing $\mu$ in Eq. (8) with $v_{\text{field}}$, keeping the observed velocities $v_{r}$ and their corresponding errors $\epsilon_{i}$ the same as in the previous numerical maximization. Instead of using data in Voronoi cells on the sky as we did previously, we now fit the entire data set. Because of the geometry of the sphere, we require that $b_{\text{apex}}$ lies in the range $-90^{\circ}$ to $90^{\circ}$ as allowing the latitude to extend past these values is unphysical, but impose no bounds on the other parameters to remain agnostic. The parameter errors are again computed using Eq. (9).111111 The code used for performing the fit of Eq. (10) and a tutorial on how to use it is given in https://github.com/abystrom/apex-likelihood-fitting.
@@
The parameters presented in Table 1 are similar to those in the halo velocity model by PP21, which is also used by Y24 and Chandra et al. (2024): our $v_{\text{compr}}$ corresponds to their mean halo radial velocity $\langle v_{r}\rangle$; $v_{\text{dipole}}$ to their MW disc barycentre velocity $v_{\text{travel}}$; and our $(l_{\text{apex}},b_{\text{apex}})$ to theirs. The model they use constrains the rotation of the halo by including transverse velocities, while we only use radial velocities in our velocity field parametrization. These radial velocities are measured from the position of the Sun, and due to the Sun not being at the centre of the Galaxy, our radial velocities will show the dipole and some contribution from the halo’s rotation projected onto our line-of-sight. However we show in Appendix E that for Galactocentric distances larger than 20 kpc the rotation of the halo has a negligible contribution to these radial velocities, and so we can compare their parameters to ours [^12] PP21 and Y24 use several different tracer populations, but we only compare to the results using their BHB sample.. We see that the velocities as well as $b_{\text{apex}}$ are consistent between all papers using BHB stars as tracers. However, our $l_{\text{apex}}$ does not agree with either work; this means that our velocity field has a different direction than their studies. The resulting angular offsets between the apex directions are $76^{\circ}$ and $57^{\circ}$, respectively. This substantial difference means that the velocity field models predict different velocities across the sky. Our apex direction ($l_{\text{apex}},b_{\text{apex}}$) is seen on-sky in Fig. 9 in blue together with the BHB sample in light grey, and the apex directions found by Y24, PP21 and Chandra et al. (2024) in pink, magenta and purple, respectively, with the position of the LMC as an orange star with its past orbit also in orange. Note that while this work and Y24 use $R>50$ kpc, the works by PP21, Chandra et al. (2024) use $R>40$ kpc.
@@
Figure 9: The apex direction longitude and latitude ($l_{\text{apex}},b_{\text{apex}}$) from Table 1 shown on-sky in blue, and the apex directions from Chandra et al. (2024), PP21 and Y24 as a purple square, magenta upwards triangle and a violet downwards triangle respectively. For PP21 we use their BHB-derived value (using Galactocentric distances larger than 40 kpc), and for Y24, we use their BHB-derived apex direction for the distance bin with Galactocentric distances larger than 50 kpc. The distribution of the BHB sample is shown in light gray. The position of the LMC is shown as an orange star with its past trajectory in orange (Vasiliev, 2023).